A hall has a volume of 240m3. An air conditioning system is to replace the air im this hall every thirty minutes using ducts that have a square cross section. Assuming that air can be treated as an incrompessible fluid, find the length of a side of the square if the air speed within the ducts is 6.0 m/s.


1
Expert's answer
2021-05-30T13:28:00-0400
"V=a^2vt\\240=(6)(30\cdot60)a^2\\a=0.15\ m"